OpenStudy (anonymous):

Find the roots of the quadratic equation x^2-5x-24=0

OpenStudy (experimentx):

x=8,-3

OpenStudy (anonymous):

can you explain how you did it please

OpenStudy (experimentx):

x^2-8x+3x-24 =0 x(x-8)+3(x-8)=0 (x-8)(x+3)=0 either x-8=0 or x+3=0 => x=8 or x=-3

OpenStudy (anonymous):

thanks